OpenStudy (anonymous):

Show that w= [3,-5] can be written as a linear combination of the vectors [2,7] [1,-1] [-2,1] (All vertical vectors btw)

ganeshie8 (ganeshie8):

like this : \[ W= \left[ {\begin{array}{cc} 3 \\ -5\\ \end{array} } \right] , A = \left[ {\begin{array}{cc} 2 \\ 7\\ \end{array} } \right], B = \left[ {\begin{array}{cc} 1 \\ -1\\ \end{array} } \right], C = \left[ {\begin{array}{cc} -2\\ 1\\ \end{array} } \right] \]

ganeshie8 (ganeshie8):

you wanto write \( W\) as linearcombination of \(A, B, C\), right ?

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

lets see.. guess we need to find it trial and error

OpenStudy (anonymous):

there is way to do it by just putting it int a matrix but the thing is there is 2 equations and 3 unknowns so that is why I am confused

ganeshie8 (ganeshie8):

W = Ax + By + Cz you need to find x, y, z

ganeshie8 (ganeshie8):

so yes, two equations, and 3 unknowns

ganeshie8 (ganeshie8):

3 = 2x+y-2z --------(1) -5 = 7x-y+z---------(2)

ganeshie8 (ganeshie8):

you will get infinite solutions, just pick one ?

OpenStudy (anonymous):

oh okay how do you know it will be infinite soln though?

ganeshie8 (ganeshie8):

look at each of the equation u have, it has 3 variables eh ?

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

that means, its a plane in 3 dimensions. 2 planes meet in a line

ganeshie8 (ganeshie8):

a line has infinite points. so infinite solutions

ganeshie8 (ganeshie8):

lets find them..

ganeshie8 (ganeshie8):

few of them atleast :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

:)

ganeshie8 (ganeshie8):

3 = 2x+y-2z --------(1) -5 = 7x-y+z---------(2) (1) + (2) gives us, -2 = 9x-z ----------(3)

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

thats ur solution line, every point on that line (3) is a solution. which means, u can use any point on that line, that gives u linear combination of A B C which equals W

ganeshie8 (ganeshie8):

say, x =1, put this in (3), and solve z -2 = 9(1) - z z = 11

ganeshie8 (ganeshie8):

put x=1, z=11 in (1) and solve y 3 = 2(1) +y-2(11) y = 23

ganeshie8 (ganeshie8):

so, when x =1, we have y=23, z=11

ganeshie8 (ganeshie8):

check if this combination works

ganeshie8 (ganeshie8):

W = Ax+By+Cz

OpenStudy (anonymous):

it workss

ganeshie8 (ganeshie8):

we could have taken x=0, that would have simplified the calculaiton, but its ok.. u see that all points on that line can be used right ?

OpenStudy (anonymous):

im still ed how adding 1 and 2 gives you the line of intersection though?

OpenStudy (anonymous):

confused*

ganeshie8 (ganeshie8):

ohk, thats simple, i just got rid of one variable by 'elimination method'

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

so, if i understand ur question correctly, you are asking, why adding both the equations gives line of intersection of planes ?

OpenStudy (anonymous):

yeah :)

ganeshie8 (ganeshie8):

thats really very good question:) before i answer that, let me ask you a similar q, u knw that intersection of two lines is a point, right ?

ganeshie8 (ganeshie8):

for ex, take below two lines :- x+y = 2 x+2y = 1

OpenStudy (anonymous):

yep and the intersection of two panes is a line

ganeshie8 (ganeshie8):

since they're not parallel, they will meet at a point for sure.

ganeshie8 (ganeshie8):

now tell me, how u wud find that intersection point ?

ganeshie8 (ganeshie8):

tell me how to find intersecting point of below two lines, x+y = 2 x+2y = 1

OpenStudy (anonymous):

equal them to eachother

ganeshie8 (ganeshie8):

yes, why it works ?

ganeshie8 (ganeshie8):

by adding the two planes vertically, I did the same previously

OpenStudy (anonymous):

oh

ganeshie8 (ganeshie8):

you're trying to figure out, why adding them vertically gives the line of intersection ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

we can find answer to that, once we find answer to below :- find intersection point of below two lines x+y = 2 x+2y = 1

ganeshie8 (ganeshie8):

follow me closely if u can, to find intersection point, lets take the first equation x+y = 2

ganeshie8 (ganeshie8):

we can add/subtract same thing to both sides, right ? so lets add "x+2y" to both sides

ganeshie8 (ganeshie8):

*subtract actually.. x+y = 2 subtract "x+2y" from both sides x+y - (x+2y) = 2 -(x+2y)

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

simplify left side

ganeshie8 (ganeshie8):

notice that, till now, 2nd equation dint come into picture

ganeshie8 (ganeshie8):

x+y = 2 subtract "x+2y" from both sides x+y - (x+2y) = 2 -(x+2y) x+y -x-2y = 2-(x+2y) -y = 2-(x+2y)

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

Now, look at 2nd equation, x+2y=1, so put that value on right side. (this step actually determines the solution we get lies on second equation also)

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

x+y = 2 subtract "x+2y" from both sides x+y - (x+2y) = 2 -(x+2y) x+y -x-2y = 2-(x+2y) -y = 2-(x+2y) from 2nd equation, x+2y=1 -y = 2-1 y = -1

ganeshie8 (ganeshie8):

since you knw y, u can find x x = 3 so, (3, -1) lies on both sides

OpenStudy (anonymous):

oohhh ok

ganeshie8 (ganeshie8):

if u are convinced (3,-1) lies on both lines, then u will understand elimination method. in elimination method also, when we add equations vertically, exact same thing happens

ganeshie8 (ganeshie8):

so, you started linear algebra is it

OpenStudy (anonymous):

yep.

ganeshie8 (ganeshie8):

i did this few months back gilbert strang's course... it was very good, if u get extra time, go thru lectures... they're very helpful :)

OpenStudy (anonymous):

ooh okay thanks so much for your help!

ganeshie8 (ganeshie8):

np :)